OpenStudy (anonymous):

Lane is 10 times Mel's age. If the difference in their age is 27, how old is Mel?

OpenStudy (perl):

L = 10*M L - M = 27 Can you solve now?

OpenStudy (anonymous):

tres

OpenStudy (anonymous):

:))))))

OpenStudy (anonymous):

You need to just solve the above two equations that perl has made them for you.. :)

OpenStudy (anonymous):

Thank you guys!

OpenStudy (anonymous):

Here Mel age is 2.7 and Lane age is 29.7 Respectively

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[\{L=10 M,L-M=27\} \]\[\{L=30,M=3\} \]